OpenStudy (anonymous):

What is the product of 4 the square root of 6 • 6 the square root of 2? Simplify if possible.

OpenStudy (anonymous):

24the square root of 2. 48the square root of 2. 24the square root of 3. 48the square root of 3.

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large 4\sqrt{6}*6\sqrt{2}\] \[\Large 4*6*\sqrt{6*2}\] \[\Large 24*\sqrt{12}\] and let you finish up

OpenStudy (anonymous):

24 the square root of 3

jimthompson5910 (jim_thompson5910):

how would you simplify \[\Large \sqrt{12}\]

OpenStudy (kamibug):

The simplified form of the sqrt of 12 is 2sqrt3. So your final answer should be 48sqrt3 :) Hope this helps!